# ARCTraj: A Dataset and Benchmark of Human Reasoning Trajectories for Abstract Problem Solving

### v2

@@ C.3. Shared Intentions Across Participants
Figure 9. State transition in ARCLE (lee2024arcle): the agent observes the current grid and applies a symbolic editing action to update it. This process models ARC as a Markov decision problem, where the agent sequentially edits the grid to reach a correct solution.:
@@ D.1. ARCLE and RL Framework
ARCLE (lee2024arcle) formulates the ARC challenge as a reinforcement learning (RL) problem, where the output grid acts as the environment and symbolic editing operations define the discrete action space. This framing enables end-to-end learning of sequential decision-making policies directly over symbolic grid transformations.
@@
ARCLE uses ARCTraj’s actionSequence to learn a mapping from states to expert actions, and then fine-tunes the policy with PPO to generalize beyond demonstrations and handle unseen states. Follow-up work (lee2024analogical) integrates DreamerV3 with ARCTraj-based initialization, further improving sample efficiency and planning.
@@ D.2. World Model Training with ARCTraj Data
Building on the RL framework, ARCLE incorporates a DreamerV3-based World Model (hafner2023dreamerv3) to learn compact latent representations of ARC tasks from the richly annotated ARCTraj data. This latent modeling improves planning and policy learning by mapping high-dimensional grid and symbolic action states into a continuous, lower-dimensional space that preserves essential task features.
@@
Figure 10. The DreamerV3-based World Model (lee2024analogical) is trained on ARCTraj using an encoder–decoder architecture that embeds grid states and symbolic operations into latent variables $z_{t}$. A recurrent state $h_{t}$ captures temporal dependencies between latents and actions $a_{t}$. Minimizing reconstruction loss enables accurate recovery of states and operations. Using ARCTraj’s rich trajectories, the model predicts future states and actions in a compact latent space, improving planning, sample efficiency, and generalization on ARC tasks.:
@@
Figure 11. LDCQ trains an autoencoder on ARCTraj’s symbolic action sequences to learn a continuous latent space for discrete operations (kim2024diffusion). The encoder embeds action trajectories into latent representations, and the decoder reconstructs actions from them. Constrained Q-learning then optimizes a policy within this latent space, using rewards to guide exploration and ensure valid solutions. This continuous space enables smoother diffusion-based sampling, thereby reducing the challenges associated with large discrete action spaces. Grounding the latent space in human trajectories from ARCTraj further improves sample efficiency and solution quality.:
@@ D.3. Diffusion-based Offline RL
LDCQ (Latent Diffusion with Constrained Q-Learning) (kim2024diffusion) introduces an approach that integrates latent diffusion models with constrained Q-learning to address the compositional and multi-step nature of ARC tasks. Instead of operating directly in a discrete symbolic action space, where optimization becomes difficult due to sparsity and discontinuity, LDCQ constructs a continuous latent action space using an autoencoder trained on ARCTraj’s rich symbolic trajectories. This latent space captures semantic relationships among symbolic operations, placing functionally similar actions closer together and enabling smoother transitions during policy learning. By learning such structure-aware embeddings, the model gains the ability to reason over symbolic operations more flexibly and expressively.
@@ D.4. GFlowNet-based Trajectory Augmentation
The GFlowNet-based approach (hwang2025gfn) frames ARC task solving as a structured sequence generation problem, where each trajectory is a series of symbolic editing operations that transform the input grid into the target output. Instead of optimizing for a single best solution as in traditional reinforcement learning, GFlowNet learns a generative policy that samples trajectories with probabilities proportional to their rewards. This probabilistic formulation facilitates the discovery of diverse, high-quality solutions that more accurately reflect the variety of human problem-solving strategies.
@@
Figure 12. Overview of the GFlowNet-based ARC solver that leverages ARCTraj’s human trajectories as reward-supporting paths (hwang2025gfn). The model generates symbolic editing sequences with probabilities proportional to their rewards, and the human trajectories guide flow matching toward high-quality solution paths. This probabilistic training promotes diverse exploration and efficient learning across ARC’s complex, multi-modal solution landscape.:
@@ D.5. Decision Transformer
Recent work (park2023unraveling) frames ARC task solving as a sequential decision-making problem and applies Decision Transformers (DTs) to model user behavior from offline data. DTs treat the task as sequence modeling, predicting the next action conditioned on the current grid state, past actions, and the expected future returns (return-to-go). By converting ARCTraj’s symbolic trajectories into (return-to-go, state, action) triples, DTs learn human problem-solving strategies directly from offline logs. Our work extends this by adding object-level information, forming (return-to-go, state, action, object) quadruplets to better capture spatial reasoning.
@@
Figure 13. Decision Transformer-based ARC solver trained on ARCTraj’s symbolic action sequences and return-to-go signals (park2023unraveling). The transformer receives sequences of past returns, states, actions, and optionally object or intention embeddings, enabling it to predict next editing operations and replicate human-like reasoning. This offline framework leverages rich human trajectories to improve policy learning without costly environment rollouts.:
Building on this framework, recent work (kim2025intention) incorporates high-level intention embeddings from user logs, forming pentaplet samples (return-to-go, state, action, object, intention). This additional conditioning enables the DT to model more abstract planning behaviors and produce coherent, higher-level action sequences that generalize better to unseen ARC tasks.
@@ E.1. Comparison with H-ARC Dataset
We compare ARCTraj with H-ARC (legris2024harc), a dataset of human editing sequences on ARC tasks. H-ARC logs pixel-level edits, where each action merges selection and color change. ARCTraj instead captures object-level sequences via the O2ARC interface (shim2024o2arc), which separates selection from symbolic operations such as Move, Rotate, Flip, Copy, and Paste. This object-centric design reveals higher-level structural intentions that pixel-level logs cannot capture.

### v3

@@ C.3. Shared Intentions Across Participants
Figure 9. State transition in ARCLE (Lee et al., 2024a): the agent observes the current grid and applies a symbolic editing action to update it. This process models ARC as a Markov decision problem, where the agent sequentially edits the grid to reach a correct solution.
State transition in ARCLE: the agent observes the current grid and applies a symbolic editing action to update it. This process models ARC as a Markov decision problem, where the agent sequentially edits the grid to reach a correct solution.
@@ D.1. ARCLE and RL Framework
ARCLE (Lee et al., 2024a) formulates the ARC challenge as a reinforcement learning (RL) problem, where the output grid acts as the environment and symbolic editing operations define the discrete action space. This framing enables end-to-end learning of sequential decision-making policies directly over symbolic grid transformations.
@@
ARCLE uses ARCTraj’s actionSequence to learn a mapping from states to expert actions, and then fine-tunes the policy with PPO to generalize beyond demonstrations and handle unseen states. Follow-up work (Lee et al., 2024b) integrates DreamerV3 with ARCTraj-based initialization, further improving sample efficiency and planning.
@@ D.2. World Model Training with ARCTraj Data
Building on the RL framework, ARCLE incorporates a DreamerV3-based World Model (Hafner et al., 2023) to learn compact latent representations of ARC tasks from the richly annotated ARCTraj data. This latent modeling improves planning and policy learning by mapping high-dimensional grid and symbolic action states into a continuous, lower-dimensional space that preserves essential task features.
@@
Figure 10. The DreamerV3-based World Model (Lee et al., 2024b) is trained on ARCTraj using an encoder–decoder architecture that embeds grid states and symbolic operations into latent variables $z_{t}$. A recurrent state $h_{t}$ captures temporal dependencies between latents and actions $a_{t}$. Minimizing reconstruction loss enables accurate recovery of states and operations. Using ARCTraj’s rich trajectories, the model predicts future states and actions in a compact latent space, improving planning, sample efficiency, and generalization on ARC tasks.
[…]
@@
Figure 11. LDCQ trains an autoencoder on ARCTraj’s symbolic action sequences to learn a continuous latent space for discrete operations (Kim et al., 2024). The encoder embeds action trajectories into latent representations, and the decoder reconstructs actions from them. Constrained Q-learning then optimizes a policy within this latent space, using rewards to guide exploration and ensure valid solutions. This continuous space enables smoother diffusion-based sampling, thereby reducing the challenges associated with large discrete action spaces. Grounding the latent space in human trajectories from ARCTraj further improves sample efficiency and solution quality.
[…]
@@ D.3. Diffusion-based Offline RL
LDCQ (Latent Diffusion with Constrained Q-Learning) (Kim et al., 2024) introduces an approach that integrates latent diffusion models with constrained Q-learning to address the compositional and multi-step nature of ARC tasks. Instead of operating directly in a discrete symbolic action space, where optimization becomes difficult due to sparsity and discontinuity, LDCQ constructs a continuous latent action space using an autoencoder trained on ARCTraj’s rich symbolic trajectories. This latent space captures semantic relationships among symbolic operations, placing functionally similar actions closer together and enabling smoother transitions during policy learning. By learning such structure-aware embeddings, the model gains the ability to reason over symbolic operations more flexibly and expressively.
@@ D.4. GFlowNet-based Trajectory Augmentation
The GFlowNet-based approach (Hwang et al., 2024) frames ARC task solving as a structured sequence generation problem, where each trajectory is a series of symbolic editing operations that transform the input grid into the target output. Instead of optimizing for a single best solution as in traditional reinforcement learning, GFlowNet learns a generative policy that samples trajectories with probabilities proportional to their rewards. This probabilistic formulation facilitates the discovery of diverse, high-quality solutions that more accurately reflect the variety of human problem-solving strategies.
@@
Figure 12. Overview of the GFlowNet-based ARC solver that leverages ARCTraj’s human trajectories as reward-supporting paths (Hwang et al., 2024). The model generates symbolic editing sequences with probabilities proportional to their rewards, and the human trajectories guide flow matching toward high-quality solution paths. This probabilistic training promotes diverse exploration and efficient learning across ARC’s complex, multi-modal solution landscape.
[…]
@@ D.5. Decision Transformer
Recent work (Park et al., 2023) frames ARC task solving as a sequential decision-making problem and applies Decision Transformers (DTs) to model user behavior from offline data. DTs treat the task as sequence modeling, predicting the next action conditioned on the current grid state, past actions, and the expected future returns (return-to-go). By converting ARCTraj’s symbolic trajectories into (return-to-go, state, action) triples, DTs learn human problem-solving strategies directly from offline logs. Our work extends this by adding object-level information, forming (return-to-go, state, action, object) quadruplets to better capture spatial reasoning.
@@
Figure 13. Decision Transformer-based ARC solver trained on ARCTraj’s symbolic action sequences and return-to-go signals (Park et al., 2023). The transformer receives sequences of past returns, states, actions, and optionally object or intention embeddings, enabling it to predict next editing operations and replicate human-like reasoning. This offline framework leverages rich human trajectories to improve policy learning without costly environment rollouts.
[…]
Building on this framework, recent work (Kim et al., 2025) incorporates high-level intention embeddings from user logs, forming pentaplet samples (return-to-go, state, action, object, intention). This additional conditioning enables the DT to model more abstract planning behaviors and produce coherent, higher-level action sequences that generalize better to unseen ARC tasks.
@@ E.1. Comparison with H-ARC Dataset
We compare ARCTraj with H-ARC (LeGris et al., 2025), a dataset of human editing sequences on ARC tasks. H-ARC logs pixel-level edits, where each action merges selection and color change. ARCTraj instead captures object-level sequences via the O2ARC interface (Shim et al., 2024), which separates selection from symbolic operations such as Move, Rotate, Flip, Copy, and Paste. This object-centric design reveals higher-level structural intentions that pixel-level logs cannot capture.